An airplane flew 4 hours with a 25 mph tail wind. The return trip against the same wind took 5 hours. Find the speed of the airplane in still air. : An airplane flew 4 hours with a 25 mph tail wind. The return trip against the same wind took 5 hours. Find the speed of the airplane in still air.
Answer by josmiceli(2158)

s = speed in still air
(s+25)*4 = (s-25)*5
4s + 100 = 5s - 125
s = 225 mph